# references.bib

### v1

@@ Abstract
Figure 1: Performance comparision of GPT-4o Audio1, Kimi-Audio [ding2025kimi], Qwen-Omni2 and Step-Audio 2 and on various benchmarks.
@@ 1 Introduction
11footnotetext: GPT-4o Audio is evaluated with gpt-4o-transcribe for ASR and gpt-4o-audio-preview-2025-06-03 for others via official API.22footnotetext: Qwen-Omni is evaluated with Qwen2.5-Omni for ASR, MMAU and speech-to-text translation, and qwen-omni-turbo-2025-03-26 for others via official API.
@@
To ensure its intelligence and expressiveness in diverse conversational scenarios, we carefully design a multi-stage training strategy to train Step-Audio 2 on 680 billion tokens of text data and 8 million hours of real and synthesized audio data. Evaluation results shown in Figure 1 demonstrate that Step-Audio 2 achieves state-of-the-art performance in a series of audio tasks, including automatic speech recognition (ASR) on multiple languages, audio understanding, speech-to-speech translation (S2ST) and speech-to-speech conversation.
@@ 4.1 Automatic speech recognition
As the most critical component of audio understanding and speech interaction, we first evaluate the model’s capability in automatic speech recognition. We evaluate Step-Audio 2 across six Chinese test sets, four English test sets, three multilingual test sets (Japanese, Cantonese, Arabic), and six in-house Chinese dialect and accented Mandarin test sets. For comparative analysis, we utilize top-performing models from both open-source and commercial domains as baselines, including Doubao LLM ASR [^1] Doubao LLM ASR refers to https://www.volcengine.com/docs/6561/1354868, GPT-4o Transcribe [^2] GPT-4o Transcribe is evaluated using its latest model, gpt-4o-transcribe, via its official API., Kimi-Audio [ding2025kimi], and Qwen2.5-Omni [xu2025qwen25omnitechnicalreport]. We prefer GPT-4o Transcribe than GPT-4o Audio, Qwen2.5-Omni than Qwen-Omni since the formers provide stronger results. Notably, Doubao LLM ASR and GPT-4o Transcribe represent specialized ASR systems that achieve leading-edge performance.
As the results shown in Table 1, Step-Audio 2 outperforms existing open-source and commercial ASR models in both general English and Chinese recognition, achieving an average word error rate (WER) of 3.18% on English and an average character error rate (CER) of 3.11% on Chinese test sets. Moreover, Step-Audio 2 offers comparable results to GPT-4o Transcribe on Arabian and Japanese recognition, to Qwen2.5-Omni on Cantonese recognition, demonstrating its capability in multilingual speech recognition. In addition, Step-Audio 2 achieves the lowest average CER among 4 in-house Chinese accented Mandarin and 2 dialect test sets. These results highlight the superiority of Step-Audio 2 in understanding the semantic information in speech.
Table 1: Comparison between Doubao LLM ASR, GPT-4o Transcribe, Kimi-Audio, Qwen-Omni and Step-Audio 2, on character (for Chinese, Cantonese and Japanese) and word (for Arabian and English) error rates among multiple ASR test sets.
Category Test set Doubao LLM ASR GPT-4o Transcribe Kimi- Audio Qwen2.5- Omni Step- Audio 2
@@ 4.2 Paralinguistic information understanding
We then evaluate how Step-Audio 2 understands the paralinguistic information in speech beyond the semantic information. To this end, we introduce Step-Audio Paralinguistic, a speech-to-speech benchmark that evaluates the model’s understanding of paralinguistic information across 11 dimensions using single-turn question answering.
Step-Audio Paralinguistic comprises 550 speech samples evenly distributed across 11 tasks. We initially collect 400 Chinese speech clips for 8 of these tasks from public podcast recordings, encompassing gender, age, timbre, emotion, pitch, rhythm, speaking speed, speaking style, and vocal activity prediction or description. For sound event, scenario, and vocal sound detection or description, we source 50 event-related, 50 environmental, and 50 vocal sounds from AudioSet [7952261], CochlScene [9979822], and VocalSound [9746828], respectively. All original recordings are shorter than 30 seconds and uniformly resampled to 24,000 Hz, with annotations provided by professional groups in open-set natural language.
@@
We also establish an automatic evaluation protocol for Step-Audio Paralinguistic, which initially transcribes model outputs into text using ASR, followed by automatic judgment with a textual LLM. More information, along with the complete Step-Audio Paralinguistic test set and evaluation code, is available at https://github.com/stepfun-ai/Step-Audio2 to foster further research on paralinguistic information understanding.
We evaluate GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, and Step-Audio 2 using the Step-Audio Paralinguistic benchmark, with results presented in Table 2. The experimental results highlight the comprehensive capabilities of Step-Audio 2 in understanding various paralinguistic information, achieving an average accuracy of 76.55, which is a significant improvement over other baseline models.
Table 2: Comparison between GPT4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA and Step-Audio 2 on the Step-Audio Paralinguistic test set.
@@ 4.3 Audio understanding
We then assess Step-Audio 2’s general audio comprehension across sound, speech, and music using the MMAU-v05.15.25 benchmark [sakshi2024mmau].
As baselines, we employ Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1 [rouditchenko2025omni], Qwen2.5-Omni, and Step-Audio-AQAA. We obtain the reported results for Audio Flamingo 3, Omni-R1, and Qwen2.5-Omni from their original papers. The results of Gemini 2.5 Pro are obtained from the official website of MMAU. And we re-evaluate GPT-4o Audio, Kimi-Audio, Step-Audio-AQAA due to the recent update of the MMAU benchmark.
@@
Table 3: Comparison between Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1, Qwen2.5-Omni, Step-Audio-AQAA and Step-Audio 2 on the MMAU-v05.15.25 Test-mini.
@@ 4.5 Tool calling
To address the gap in the availability of suitable test sets for tool calling in speech conversations, we introduce Step-Audio Toolcall, a test set that evaluates the model’s ability in tool invocation, selection and parameter extraction under Chinese speech conversations.
We employ a textual LLM to generate 200 multi-turn dialogue scripts for each kind of tool. Each script contains 3-6 turns of inputs and outputs, in which previous turns may or may not include tool calling statements, but the final input must contain a calling intention to a specific external tool. We then balance the samples with an equal number of negative samples for each kind of tools, in which the final speech input either has no tool calling intention or intention to call on other kinds of tools. Subsequently, we synthesize these scripts into speeches with our conversation synthesis pipeline. And we propose an automatic evaluation protocol to employ Qwen3-32B to automatically examine the output and tool calling statements. We release Step-Audio Toolcall including the original scripts, synthesized speech conversations and the corresponding evaluation script in https://github.com/stepfun-ai/Step-Audio2.
@@
Table 5: Comparison between Step-Audio 2 and Qwen3-32B on tool calling. Qwen3-32B is evaluated with text inputs. Date and time tools have no parameter.
@@ 5 Conclusion
We introduce Step-Audio 2, an end-to-end large audio language model (LALM) designed for enterprise speech and audio understanding, as well as intelligent speech interaction. Step-Audio 2 leverages a latent audio encoder and reinforcement learning to enhance its speech and audio comprehension capabilities. Furthermore, by integrating the generation of discrete audio tokens into language modeling, Step-Audio 2 achieves genuine end-to-end speech interaction and improves its responsiveness to paralinguistic information, such as speaking styles and emotions. Step-Audio 2 is also capable of utilizing external tools including web search and audio search for multi-modal RAG. Trained on 8 million hours of speeches and audios, Step-Audio 2 demonstrates state-of-the-art performance across various tasks, including ASR, audio understanding, speech translation, and general speech conversation, outperforming both open-source and commercial solutions.

### v2

@@ Abstract
Figure 1: Performance comparision of GPT-4o Audio1, Kimi-Audio2, Qwen-Omni3 and Step-Audio 2 and on various benchmarks.
@@ 1 Introduction
11footnotetext: GPT-4o Audio is evaluated with gpt-4o-transcribe for ASR and gpt-4o-audio-preview-2025-06-03 for others via official API.22footnotetext: Kimi-Audio is excluded from translation evaluations since it consistently ignores prompts.33footnotetext: Qwen-Omni is evaluated with Qwen2.5-Omni for MMAU and speech-to-text translation, and qwen-omni-turbo-2025-03-26 for others via official API.
@@
To ensure its intelligence and expressiveness in diverse conversational scenarios, we carefully design a multi-stage training strategy to train Step-Audio 2 on 680 billion tokens of text data and 8 million hours of real and synthesized audio data. Evaluation results shown in Figure 1 demonstrate that Step-Audio 2 achieves state-of-the-art performance in a series of audio tasks, including automatic speech recognition (ASR) on multiple languages, audio understanding, speech-to-speech translation and speech-to-speech conversation. Typical usages of Step-Audio 2 are also illustrated in Figure 2.
Figure 2: Illustration of the applications of Step-Audio 2 across various speech conversation scenarios.
@@ 4.1 Automatic speech recognition
As the most critical component of audio understanding and speech interaction, we first evaluate the model’s capability in automatic speech recognition. We evaluate Step-Audio 2 across six Chinese test sets, four English test sets, three multilingual test sets (Japanese, Cantonese, Arabic), and six in-house Chinese dialect and accented Mandarin test sets. For comparative analysis, we utilize top-performing models from both open-source and commercial domains as baselines, including Doubao LLM ASR [^1] Doubao LLM ASR refers to https://www.volcengine.com/docs/6561/1354868, GPT-4o Transcribe [^2] GPT-4o Transcribe is evaluated using its latest model, gpt-4o-transcribe, via its official API., Kimi-Audio [ding2025kimi], and Qwen-Omni. We prefer GPT-4o Transcribe than GPT-4o Audio since the formers provide stronger results. Notably, Doubao LLM ASR and GPT-4o Transcribe represent specialized ASR systems that achieve leading-edge performance.
We evaluate all the models without specifying language [^3] We evaluate without specifying language to ensure a fair comparison. Notably, Qwen-Omni lacks a language-independent testing approach, specifying language may yield better results. and summarize the results in Table 1. Step-Audio 2 outperforms existing open-source and commercial ASR models in both general English and Chinese recognition, achieving an average word error rate (WER) of 3.18% on English and an average character error rate (CER) of 3.11% on Chinese test sets. Moreover, Step-Audio 2 offers comparable results to GPT-4o Transcribe on Arabian and Japanese recognition, to Qwen-Omni on Cantonese recognition, demonstrating its capability in multilingual speech recognition. In addition, Step-Audio 2 achieves the lowest average CER among 4 in-house Chinese accented Mandarin and 2 dialect test sets. These results highlight the superiority of Step-Audio 2 in understanding the semantic information in speech.
Table 1: Comparison between Doubao LLM ASR, GPT-4o Transcribe, Kimi-Audio, Qwen-Omni and Step-Audio 2, on character (for Chinese, Cantonese and Japanese) and word (for Arabian and English) error rates among multiple ASR test sets. N/A indicates that the language is not supported.
Category Test set Doubao LLM ASR GPT-4o Transcribe Kimi- Audio Qwen- Omni Step- Audio 2
@@ 4.2 Paralinguistic information understanding
We then evaluate how Step-Audio 2 understands the paralinguistic information in speech beyond the semantic information. To this end, we introduce StepEval-Audio-Paralinguistic, a speech-to-speech benchmark that evaluates the model’s understanding of paralinguistic information across 11 dimensions using single-turn question answering.
StepEval-Audio-Paralinguistic comprises 550 speech samples evenly distributed across 11 tasks. We initially collect 400 Chinese speech clips for 8 of these tasks from public podcast recordings, encompassing gender, age, timbre, emotion, pitch, rhythm, speaking speed, speaking style, and vocal activity prediction or description. For sound event, scenario, and vocal sound detection or description, we source 50 event-related, 50 environmental, and 50 vocal sounds from AudioSet [7952261], CochlScene [9979822], and VocalSound [9746828], respectively. All original recordings are shorter than 30 seconds and uniformly resampled to 24,000 Hz, with annotations provided by professional groups in open-set natural language.
@@
We also establish an automatic evaluation protocol for StepEval-Audio-Paralinguistic, which initially transcribes model outputs into text using ASR, followed by automatic judgment with a textual LLM. More information, along with the complete StepEval-Audio-Paralinguistic test set and evaluation code, is available at https://github.com/stepfun-ai/Step-Audio2 to foster further research on paralinguistic information understanding.
We evaluate GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, and Step-Audio 2 using the StepEval-Audio-Paralinguistic benchmark, with results presented in Table 2. The experimental results highlight the comprehensive capabilities of Step-Audio 2 in understanding various paralinguistic information, achieving an average accuracy of 76.55, which is a significant improvement over other baseline models.
Table 2: Comparison between GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA and Step-Audio 2 on StepEval-Audio-Paralinguistic.
@@ 4.3 Audio understanding
We then assess Step-Audio 2’s general audio comprehension across sound, speech, and music using the latest version of the MMAU benchmark [sakshi2024mmau] [^4] MMAU v05.15.25 test-mini.
As baselines, we employ Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1 [rouditchenko2025omni], Qwen2.5-Omni, and Step-Audio-AQAA. We obtain the reported results for Audio Flamingo 3, Omni-R1, and Qwen2.5-Omni from their original papers. The results of Gemini 2.5 Pro are obtained from the official website of MMAU. And we re-evaluate GPT-4o Audio, Kimi-Audio and Step-Audio-AQAA due to the recent update of the MMAU benchmark.
@@
Table 3: Comparison between Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1, Qwen2.5-Omni, Step-Audio-AQAA and Step-Audio 2 on MMAU.
@@ 4.5 Tool calling
To address the gap in the availability of suitable test sets for tool calling in speech conversations, we introduce StepEval-Audio-Toolcall, a test set that evaluates the model’s ability in tool invocation, selection and parameter extraction under Chinese speech conversations.
We employ a textual LLM to generate 200 multi-turn dialogue scripts for each kind of tool. Each script contains 3-6 turns of inputs and outputs, in which previous turns may or may not include tool calling statements, but the final input must contain a calling intention to a specific external tool. We then balance the samples with an equal number of negative samples for each kind of tools, in which the final speech input either has no tool calling intention or intention to call on other kinds of tools. Subsequently, we synthesize these scripts into speeches with our conversation synthesis pipeline. And we propose an automatic evaluation protocol to employ Qwen3-32B to automatically examine the output and tool calling statements. We release StepEval-Audio-Toolcall including the original scripts, synthesized speech conversations and the corresponding evaluation script in https://github.com/stepfun-ai/Step-Audio2.
@@
Table 5: Comparison between Step-Audio 2 and Qwen3-32B on StepEval-Audio-Toolcall. Qwen3-32B is evaluated with text inputs. Date and time tools have no parameter.
@@ 5 Conclusion
We introduce Step-Audio 2, an end-to-end large audio language model designed for enterprise speech and audio understanding, as well as intelligent speech interaction. Step-Audio 2 leverages a latent audio encoder and reinforcement learning to enhance its speech and audio comprehension capabilities. Furthermore, by integrating the generation of discrete audio tokens into language modeling, Step-Audio 2 achieves genuine end-to-end speech interaction and improves its responsiveness to paralinguistic information, such as speaking styles and emotions. Step-Audio 2 is also capable of utilizing external tools including web search and audio search for multi-modal RAG. Trained on 8 million hours of speeches and audios, Step-Audio 2 demonstrates state-of-the-art performance across various tasks, including ASR, audio understanding, speech translation, and general speech conversation, outperforming both open-source and commercial solutions.